OpenStudy (anonymous):

Find the area of an equilateral triangle (regular 3-gon) with the given measurement. 9-inch perimeter A = sq. in.

OpenStudy (anonymous):

Im confused is 

OpenStudy (anonymous):

A = 1/2aP Note, there is also a formula for finding the area another way which is what I will use. That is A = 1/4(√3)(s^2) A = 1/4(√3)(9) A = 2.25√3 or 9/4(√3)